Net income $         12,500
Profit margin 7.80%
Accounts receivable $           6,700
Percent credit sales 85%
What is Firm's days sales in receivables?
Answers:

17.95

18.45

19.06

19.68
Answer #1

Profit margin=Net income/Sales

Sales=(12500/0.078)=$160256.41

Hence credit sales=0.85*160256.41=$136217.948

Days sales in receivables=(Accounts receivable/Credit sales)*365 days

=(6700/136217.948)*365

=17.95(Approx).